OpenStudy (anonymous):

FREE MATH TUTOR HERE Hi everyone! I will try and help you with your math questions and problems! I will tutor you over Private messages on here. Just let me know in the comments what you need help with :)

OpenStudy (anonymous):

The Maximum Error of Estimate is exactly equal to half of the width of the confidence interval. T-F, and why or why not?

OpenStudy (anonymous):

I NEED HELP!!!!!!!!! Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences. Satisfied with your answer, the wise old woman grants you the treasure of the Lost Island of Laplaya! The treasure is an awesome ability to complete the square! Way to go! Professor McMerlock is thankful he picked you and promises to call on you again for another adventure. f(x)=x^2-4. g(x)= x^2-2 h(x)=-3x^2+11i